OpenStudy (anonymous):

15b - 13 = 2 + 10b I don't want an answer, but could someone double check my thought process? (reorder the terms) 15b + 10b = -13 + 2

OpenStudy (anonymous):

@uri

OpenStudy (some.random.cool.kid):

well what is it that you have to do

OpenStudy (anonymous):

solve it. is that first step right? that is what i want to find out

OpenStudy (yanasidlinskiy):

Is your answer the last equation by..(re-order the terms) ?

OpenStudy (ccswims):

in order to solve it you must add like terms

OpenStudy (ccswims):

your on the right track but have your symbols mixed up 15b - 10b - 13 = 2

OpenStudy (anonymous):

so the after doing the adding it would look like this right? 25b = -11

OpenStudy (some.random.cool.kid):

so do you want to figure it out or do you want us to help?

OpenStudy (ccswims):

you should start off with that and work from there

OpenStudy (some.random.cool.kid):

well to get b alone

OpenStudy (anonymous):

i want you guys to check my work because i have a hard time with this stuff.

OpenStudy (some.random.cool.kid):

you have to dived to get B alone

OpenStudy (some.random.cool.kid):

at the last step

OpenStudy (anonymous):

then you would divide and it would be b = -0.44 but that isn't one of the options i have

OpenStudy (ccswims):

well, you had it right except for your signs 15b - 10b = 2 + 13

OpenStudy (ccswims):

^^ do that

OpenStudy (yanasidlinskiy):

ccswims, she is actually on the right track. If that's what her question looks like. Her symbols aren't messed up. Add Like terms on both sides. What do you get on both sides? =

OpenStudy (anonymous):

ahh ok thank you. i get those mixed up...

OpenStudy (anonymous):

so 5b = 15 which would be 3. awesome thanks guys!

OpenStudy (yanasidlinskiy):

Remember your integer rules. Don't forget. Those things throw people off:D

OpenStudy (yanasidlinskiy):

But you got it right! Congratulations!!:) Keep up the good work!

OpenStudy (anonymous):

yep. still working on that :) thank you again so much!

OpenStudy (yanasidlinskiy):

You're Most Welcome!:) Anytime!

OpenStudy (yanasidlinskiy):

Oh yes, close the question too, when answered. "Best Response" to give medals:) Just some things you might want to know:)